# MMICT: Boosting Multi-Modal Fine-Tuning with In-Context Examples

### v2

@@ 4.2 Evaluation Tasks and Datasets
For captioning tasks, we evaluate MMICT on 3 public datasets including COCO Caption (Lin et al. 2014), MSVD (Xu et al. 2017), MSR-VTT (Xu et al. 2016). We use BLEU@4 (B@4) (Papineni et al. 2002), CIDEr (C) (Vedantam, Zitnick, and Parikh 2015) as metrics. $t_{ins}$ is randomly sampled from the pre-defined instruction templates, which are shown in Tab. 1
For open-ended question answering tasks, we evaluate MMICT on 3 public datasets including VQAv2 (Goyal et al. 2017), MSVD (Xu et al. 2017) and MSR-VTT (Xu et al. 2017). We formulate them as generative problems. During inference, we use beam search with a beam width of 5 to generate answers from the whole vocabulary with no restrictions. Accuracy (Acc) is used as the evaluation metric. We evaluate VQAv2 on its validation set (the test label is not publicly available) and evaluate MSVD and MSR-VTT on their respective test sets. $t_{ins}$ for these tasks is designed as “Question: {} Answer:”.
4.3 Environment and Hyper-parameters
@@ 4.4 Baselines
  • VLAB (He et al. 2023): VLAB is a video language pre-training method that transfers CLIP’s learned representations to video-text tasks.
  • VAST (Chen et al. 2023a): VAST is an omni-modality video-text foundational model that can perceive and process vision, audio, and subtitle modalities from videos.
  • mPLUG-2 (Xu et al. 2023b): mPLUG-2 is a multi-module composition network. It contains shared modules for modality collaboration and uses different modality modules to deal with modality entanglement.
4.5 Overall Performance
@@
4.6 The Impacts of Demonstration Formats
Figure 2: Different usages of M-Hub. As demonstrated in (a) and (b), it can function as a uni-modal encoder. Moreover, it can also operate as a multi-modal fusion encoder, as shown in (c), (d), and (e).
Table 3: The results of ablation studies about ICT demonstration variants. Best results are shown in bold. For the evaluation, we train on one third data of the complate datasets that is randomly sampled.
@@ 4.6 The Impacts of Demonstration Formats
Figure 3: Case study on (a) visual question answering, (b) image captioning, (c) video question answering, and (d) video question. We show the answers generated by VanillaFT, VanillaICT-$\text{B}_{\text{VT}}$ and MMICT in orange, green and blue, respectively.
@@
Table 4: Performance using different numbers of demonstration examples.
@@
Table 5: Performance using different sample strategies for demonstration examples.
@@
4.7 The Impacts of Sampling for Demonstrations
@@
4.8 The Impacts of Demonstrations on Inference
Table 6: Results of using demonstrations during inference.
@@ 4.8 The Impacts of Demonstrations on Inference
Table 6 presents the impact of using demonstrations on the model’s performance during the inference phase. The used LLM is FlanT5. Note that the results for the default setting reported in Tab. 1 of our submission use demonstrations during fine-tuning. For each data sample in the test set, we randomly sample its demonstrations from the training set. The results lead us to the following observations:
  1. 1.
    The performance of MMICT gets slightly improved when demonstrations are incorporated during inference across most tasks. However, for the VideoQA task on the MSVD dataset, demonstrations appear to affect the performance of all methods negatively. Despite this, MMICT consistently surpasses the performance of the baselines.
  2. 2.
The above findings indicate that the use of demonstrations during inference can have varying effects on different methods and tasks. MMICT and VanillaICT-$\text{B}_{\text{VT}}$ have fully learned from demonstrations during in-context tuning and they show comparable performance regardless of whether demonstrations are leveraged during inference. Besides, the observations underscore the importance of in-context tuning in enabling MM-LLMs to learn from demonstrations effectively.
4.9 The Impacts of Feature Levels
Table 7: Performance of using different levels of features in VanillaFT.
@@ 4.9 The Impacts of Feature Levels
Tab. 7 presents the comparison of the performance of using different levels of features in VanillaFT. For frame-level features, we pass each video clip through the frozen image encoder and M-Hub individually, and then concatenate them together to obtain frame-level features. The results in Tab. 7 illustrate that, despite using 16 times fewer tokens (i.e., frame number $n_{f}$), the performance using video-level features is comparable to, and in some cases even surpasses, the performance achieved using frame-level features. The observation further demonstrates that video clips contain redundant information which does not significantly contribute to the performance. Note that, due to the limitations of the input length of LLMs, we do not conduct experiments for VanillaICT-$\text{B}_{\text{VT}}$ and MMICT, which additionally take demonstrations as input.
4.10 Direct Comparison with BLIP-2
Table 8: Results on COCO with FlanT5 as the language model. $*{}^{*}$ indicates our implementation.
Method Trainable Params COCO
B@4 C
BLIP-2 IE + Qformer 42.4 144.5
[…]
MMICT IE + M-Hub 45.7 152.6
To further demonstrate the advantages of MMICT, we provide a direct comparison with BLIP-2 in Tab. 8. As displayed in Tab. 8, the performance of MMICT shows a remarkable improvement when the image encoder is further fine-tuned, as opposed to just fine-tuning the M-Hub on the COCO dataset. Additionally, MMICT still outperforms BLIP-2 that we re-implemented under their configuration. Significantly, the performance gap between MMICT and BLIP-2 becomes even larger, indicating that the benefits of MMICT amplify as more parameters are trained.
4.11 Case Study
@@
5 Conclusion

### v3

@@ 4.2. Evaluation Tasks and Datasets
For captioning tasks, we evaluate MMICT on 3 public datasets including COCO Caption (Lin et al., 2014), MSVD (Xu et al., 2017), MSR-VTT (Xu et al., 2016). We use BLEU@4 (B@4) (Papineni et al., 2002), CIDEr (C) (Vedantam et al., 2015) as metrics. $t_{ins}$ is randomly sampled from the pre-defined instruction templates , which are shown in Tab. 1
For open-ended question answering tasks, we evaluate MMICT on 3 public datasets including VQAv2 (Goyal et al., 2017), MSVD (Xu et al., 2017) and MSR-VTT (Xu et al., 2017). We formulate them as generative problems. During inference, we use beam search with a beam width of 5 to generate answers from the whole vocabulary with no restrictions. Accuracy (Acc) is used as the evaluation metric. We evaluate VQAv2 on its validation set (the test label is not publicly available) and evaluate MSVD and MSR-VTT on their respective test sets. $t_{ins}$ for these tasks is designed as “Question: {} Answer:”.
4.3. Environment and Hyper-parameters
@@ 4.4. Baselines
  • VLAB (He et al., 2023): VLAB is a video language pre-training method that transfers CLIP’s learned representations to video-text tasks.
  • VAST (Chen et al., 2023b): VAST is an omni-modality video-text foundational model that can perceive and process vision, audio, and subtitle modalities from videos.
  • mPLUG-2 (Xu et al., 2023b): mPLUG-2 is a multi-module composition network. It contains shared modules for modality collaboration and uses different modality modules to deal with modality entanglement.
4.5. Overall Performance
@@
4.6. The Impacts of Demonstration Formats
Table 2. Performance of all methods. We mainly compare our MMICT with two baselines. The performance of SOTA methods is denoted in gray. The best results across different types of MM-LLMs are shown in bold.
Caption VQA VideoQA
COCO MSR-VTT MSVD VQAv2 MSR-VTT MSVD
Method LLM B@4 C B@4 C B@4 C Acc Acc Acc
VLAB (He et al., 2023) - 152.5 54.6 74.9 79.3 179.8 - 49.6 61.0
VAST (Chen et al., 2023b) - 149.0 - 78.0 - - - 50.1 60.2
mPLUG-2 (Xu et al., 2023b) 41.6 137.7 57.8 80.3 75.0 165.8 - 48.0 58.1
VanillaFT FlanT5 42.4 144.5 51.3 74.7 73.0 174.3 69.6 43.4 63.9
MMICT 43.6 145.7 51.6 74.8 76.5 179.1 72.3 45.6 66.7
VanillaFT OPT 43.7 145.8 47.6 69.9 80.3 177.1 56.9 42.6 56.9
MMICT 43.9 145.5 52.0 71.4 80.4 180.4 73.0 46.0 66.3
Figure 2. Different usages of M-Hub. As demonstrated in (a) and (b), it can function as a uni-modal encoder. Moreover, it can also operate as a multi-modal fusion encoder, as shown in (c), (d), and (e).
Table 3. The results of ablation studies about ICT demonstration variants. Best results are shown in bold. For the evaluation, we train on one third data of the complate datasets that is randomly sampled.
@@ 4.6. The Impacts of Demonstration Formats
Table 4. Performance using different numbers of demonstration examples.
@@
Table 5. Performance using different sample strategies for demonstration examples.
@@
4.7. The Impacts of Sampling for Demonstrations
@@ 4.7.2. Sampling Strategy
4.8. The Impacts of Demonstrations on Inference
Table 6. Results of using demonstrations during inference.
@@ 4.8. The Impacts of Demonstrations on Inference
Table 7. Results using different levels of features in VanillaFT.
@@
Tab. 6 presents the impact of using demonstrations on the model’s performance during the inference phase. The used LLM is FlanT5. Note that the results for the default setting reported in Tab. 1 of our submission use demonstrations during fine-tuning. For each data sample in the test set, we randomly sample its demonstrations from the training set. The results lead us to the following observations:
  1. (1)
    The performance of MMICT gets slightly improved when demonstrations are incorporated during inference across most tasks. However, for the VideoQA task on the MSVD dataset, demonstrations appear to affect the performance of all methods negatively. Despite this, MMICT consistently surpasses baselines.
  2. (2)
Above findings indicate that the use of demonstrations during inference can have varying effects on different methods and tasks. MMICT and VanillaICT-$\text{B}_{\text{VT}}$ have fully learned from demonstrations during in-context tuning and they show comparable performance regardless of whether demonstrations are leveraged during inference. Besides, the observations underscore the importance of in-context tuning in enabling MM-LLMs to learn from demonstrations effectively.
4.9. The Impacts of Feature Levels
Tab. 7 shows the performance of using different levels of features in VanillaFT. For frame-level features, we pass each video clip through the frozen image encoder and M-Hub individually, and then concatenate them together to obtain frame-level features. The results in Tab. 7 illustrate that, despite using 16 times fewer tokens (i.e., frame number $n_{f}$), the performance using video-level features is comparable to, and in some cases even surpasses, the performance achieved using frame-level features. The observation further demonstrates that video clips contain redundant information which does not significantly contribute to the performance. Note that, due to the limitations of the input length of LLMs, we do not conduct experiments for VanillaICT-$\text{B}_{\text{VT}}$ and MMICT, which additionally take demonstrations as input.
Figure 3. Case study on (a) visual question answering, (b) image captioning, (c) video question answering, and (d) video captioning. We show the answers generated by VanillaFT, VanillaICT-$\text{B}_{\text{VT}}$ and MMICT in orange, green and blue, respectively.
Table 8. Results on different sets of VQAv2.
Method LLM VQAv2
val test-dev
VanillaFT FlanT5 69.6 71.6
VanillaICT-$\text{B}_{\text{VT}}$ 69.8 71.8
MMICT 72.3 74.6
VanillaFT OPT 56.9 60.0
VanillaICT-$\text{B}_{\text{VT}}$ 62.7 64.0
MMICT 73.0 75.2
4.10. Case Study
@@
5. Experiments on VQAv2
Note that the test labels of VQAv2 dataset are not publicly available. Tab. 8 reports the results of MMICT compared with two baselines, i.e., VanillaFT and VanillaICT-$\text{B}_{\text{VT}}$, on the validation set and the test set. The results demonstrate that MMICT consistently outperforms the baselines on the validation set and the test set of VQAv2.
6. The differences between MMICT and Otter
Otter is the most similar wrok compared with MMICT, we have briefly discussed it in Sec. 2.2. In this section, we provide detailed comparisons between them, including the experimental results.
As shown in Tab. 9, the four main differences between MMICT and Otter are as follows:
  1. (1)
    Otter uses in-context examples during pre-training, while MMICT is a novel fine-tuning paradigm applied to downstream tasks.
  2. (2)
    The interaction of visual and textual features is implemented in M-hub. Conversely, Otter incorporates additional cross-attention modules within the layers of LLM, leading to an increase in computational requirements compared to M-hub.
  3. (3)
    The data format of Otter necessitates meticulous design, as depicted in Fig. 2 of their paper. In contrast, our data format is more flexible and does not mandate a specific design. Remarkably, our method can function effectively even with random sampling for demonstrations, as illustrated in Section 4.7.
  4. (4)
    At inference, Otter generates predictions based on the provided in-context examples. Differently, MMICT exhibits robust performance without demonstrations.
Table 9. The differences to Otter.
Method Target Stage Multi-modal Fusion In-context Unrestriction Prediction w/o Examples
Otter Pre-training LLM
MMICT Fine-tuning M-Hub
Table 10. Results compared with Otter on MSVD tasks.
Method LLM Caption VideoQA
B@4 C Acc
Otter LLaMa7B 78.9 184.2 65.2
MMICT OPT2.7B 80.4 180.4 66.3
Furthermore, as displayed in Tab. 10, we also provide a direct comparison with Otter. From the results, we can find that although the size of LLM in MMICT is much smaller than that in Otter, MMICT exists comparable and even superior performance on video captioning and VideoQA tasks when compared to Otter.
7. Conclusion
@@
Acknowledgments
This work is supported by National Science and Technology Major Project (No. 2022ZD0118201) and National Natural Science Foundation of China (No. 62002303, 42171456).